In the current year, Jill, age 35, received a job offer with two alternative compensation packages to choose from. The firstThe number needed from Exhibit12-8 is .09
Answer #1

Part a-1

Option B

Package 2 offers $82000 annual salary plus health and life insurance benefits

Jill is better off by $12490 (57910-45420) by choosing option 2

Option 1

Option 2

Salary

94000

82000

Includible health insurance

0

0

Includible life insurance

162

Includible parking

3000

Includible flight benefits

BTCF from salary

94000

85162

(1-32%)

(1-32%)

ATCF from salary

63920

57910

Cost of health insurance

(3000)

Cost of parking

(5500)

Cost of life insurance

(3000)

Cost of flight benefits

(7000)

After-tax dollars

$45420

$57910

Includible life insurance =(200000-50000)*0.09*12 = 162000/1000 = $162

Includible parking =(520-270)*12 =3000

Part a-2

After tax dollars

$12490

Part b-1

Option A

Package 1 offers $121000 annual salary with no qualified fringe benefits

Jill is better off by $5870 (63780-57910) by choosing option 1

Option 1

Option 2

Salary

121000

82000

Includible health insurance

0

0

Includible life insurance

162

Includible parking

3000

Includible flight benefits

BTCF from salary

121000

85162

(1-32%)

(1-32%)

ATCF from salary

82280

57910

Cost of health insurance

(3000)

Cost of parking

(5500)

Cost of life insurance

(3000)

Cost of flight benefits

(7000)

After-tax dollars

$63780

$57910

Includible life insurance =(200000-50000)*0.09*12 = 162000/1000 = $162

Includible parking =(520-270)*12 =3000

Part b-2

After tax dollars

$5870